OpenStudy (anonymous):

the calc company makes basic scientific calculators. each calculator costs $15.75 to produce. the fixed costs of production are 15,500 per month. these calculators sell for $35.75 each. if the company produces 55,000 calculators in february, find the cost of manufacturing them.

OpenStudy (amistre64):

15.75x+15500 = cost to manufacture

OpenStudy (amistre64):

regardless of what they sell for; the cost to produce them remains the same...

OpenStudy (amistre64):

$881,750 to produce tha tmany

OpenStudy (amistre64):

it costs about 16.04 to make just one at that rate ;)

OpenStudy (anonymous):

IF the company sells all 55,000 calculators, how mch gross income will the company generate?

OpenStudy (amistre64):

35.75 * 55000 - 881750

OpenStudy (amistre64):

(35.75 * 55,000) - 881,750 = $1,084,500

OpenStudy (amistre64):

or did i do net?

OpenStudy (amistre64):

i think i did net income :)

OpenStudy (amistre64):

35.75 * 55 000 = 1,966,250 is prolly gross

OpenStudy (anonymous):

based on the given info, hw many calculators does the company have to produce to break even and wat is the break even amount